OpenStudy (anonymous):

6. Write the equation in slope-intercept form. What are the slope and y-intercept? –12x + 11y = –8

OpenStudy (aravindg):

take a look at my tutorial http://openstudy.com/updates/50f8cc3fe4b027eb5d9a556a

OpenStudy (anonymous):

y = mx + c Slope intercept Form

OpenStudy (aravindg):

hope that helps you :)

OpenStudy (anonymous):

where m = slope c = y intercept

OpenStudy (aravindg):

@sh9down45 any doubts?

OpenStudy (anonymous):

is it y=-12/11x-8/11 slope=12/11 y=8/11

OpenStudy (anonymous):

?

OpenStudy (aravindg):

y intercept is -8/11 slope is right

OpenStudy (anonymous):

can u help me with some more?